a solution of 72% fertilizer is to be mixed with a 26% fertilizer to form 230 liters of a 53% solution. How many liters of the 72% solution must be used
Answer by stanbon(75887)

a solution of 72% fertilizer is to be mixed with a 26% fertilizer to form 230 liters of a 53% solution. How many liters of the 72% solution must be used
------
Quantity:: s + t = 230 liters
Fertilizer: 0.72s + 0.26t = 0.53*230
--------------------------------------------
Modify for elimination::
26s + 26t = 26*230
72s + 26t = 53*230
--------------------------------
Subtract and solve for "s"::
46s = 27*230
s = 135 liters (amt. of 72% solution needed)
-------------------
Cheers,
Stan H.
-------------